# Preference Poisoning Attacks on Reward Model Learning

### v1

@@ 6. Effectiveness of Defense
We study the efficacy of several state-of-the-art defense approaches in the context of poisoning attacks on reward model learning. Specifically, we consider representative defenses from three general classes of approaches (that can be applied in our setting as well) to defend against poisoning attacks: 1) identification and removal of anomalous (poisoned) data, 2) iterative training and removal of high-loss data, and 3) data randomization. As representatives of the first class, we evaluate spectral anomaly detection defense, which arises from the theoretically grounded robust learning literature (Tran et al., 2018b), and the recently proposed Meta-Sift approach (Zeng et al., 2023). In both spectral outlier and Meta-Sift defense, we first apply the respective approach to identify and remove outliers, and then train the reward model on the rest of the training data. Our representative of the second class is the following three-stage algorithm (Du et al., 2019c; Liu et al., 2017a; Vorobeychik and Kantarcioglu, 2018): 1) train an initial model $R_{\theta_{1}}$, 2) remove $\alpha\cdot B$ datapoints with the largest loss $\mathcal{L}(\tilde{D},\theta_{1})$, where $\alpha\geq 1$ is a hyperparameter (in our experiments, $\alpha=1.5$), and 3) train an updated model on the remaining data. We refer to this as the loss outlier defense. Finally, our representative of the third class of approaches is the ALIBI algorithm that leverages differential privacy and Bayesian post-processing to increase robustness to poisoning attacks (Malek Esmaeili et al., 2021).
@@
MuJoCo Control
We begin again with the MuJoCo setting, evaluating the best attack against the defensive strategies discussed above.
Figure 15. Efficacy of defense against promotion attacks in the MuJoCo setting.
Figure 16. Efficacy of defense against demotion attacks in the MuJoCo setting.
The results are shown in Figure 15 for promotion attacks and in Figure 16 for demotion attacks. In most cases, we observe that neither the spectral outlier defense nor the loss outlier defense have a significant impact on the attack efficacy. In contrast, ALIBI is considerably more efficacious in both Hopper and Walker environments, although not in the Reacher environment. Nevertheless, with 10% of the data poisoned, attacks remain relatively successful against all of the defenses.
Atari Vision-Based Control
Next, we consider the efficacy of defensive methods against the proposed attacks in the Atari vision-based control domain.
Figure 17. Efficacy of defense against promotion attack (top row) and demotion attacks (bottom row) in the Atari setting.
@@
Recommendation System
In the context of the Amazon recommendation dataset, our results of the relative defense efficacy are shown in Figure 18.
Figure 18. Efficacy of defense against promotion (left) and demotion (right) attacks in the Amazon recommendation setting.
@@
Safety Alignment
Finally, we evaluate the efficacy of defense approaches in the case of safety alignment for LLM. We exclude Meta-Sift in this setting, which requires iterative training and evaluation of the model, as it is impractical at this scale.
Figure 19. Efficacy of defense against promotion (left) and demotion (right) attacks in the safety alignment (LLM) setting. $B=$ 0.3% of data.
7. Conclusion
The problem of poisoning attacks generally, and label-flipping in particular, has received much attention in the traditional supervised learning settings, such as classification and regression. Despite some superficial similarity, however, the problem of preference poisoning is structurally distinct, both in terms of the particular optimization problem being solved in learning the reward model, and the particulars of the natural threat model involving promotion and demotion of a target set of outcomes (candidates). Our analysis yields two high-level takeaways with broad significance to security issues surrounding preference elicitation and value alignment. The first is the importance of comprehensive vulnerability analysis that involves a diverse set of attack techniques; in particular, we show that which attack among a collection is best can vary greatly by domain and environment—that is, by the dataset involved. The second is the importance of the particular dataset and data distribution. Indeed, in insightful recent work, Suya et al. (2023a) observed that the nature of the data distributions (and, thus, datasets) can be essential to robustness of models to data poisoning. This issue, which we highlight empirically, clearly deserves considerably more attention. Moreover, our observation that several state-of-the-art defense techniques fail to provide consistently strong defense in our problem setting suggests that an ability to exert control on the nature of the data we collect—something that is indeed often possible—can have substantial consequences for the vulnerabilities we have studied here.
Appendix
@@ Appendix A Additional Experiment Details
Figure 20. Demotion attack stealth in MuJoCo: test set accuracy.
Figure 21. Relative efficacy of gradient-based demotion attacks (top row) and RBD demotion attacks (bottom row) in MuJoCo.
Figure 22. Attack efficacy in MuJoCo with 5 target candidates.
Figure 23. Demotion attack stealth in Atari: test set accuracy.
Figure 24. Relative efficacy of gradient-based demotion attacks (top row) and RBD demotion attacks (bottom row) in Atari.
Figure 25. Attack efficacy in Atari with 5 target candidates.
Figure 26. Attack efficacy with 5 target candidates on the Amazon recommendation ratings dataset. Neural Network Model (top row) and Linear Model (bottom row).
@@
For the neural network structure, we explored both a simple linear model and a 3-layer neural network with ReLU activation functions and a mid-layer size of 1024. In the PCA scenario, we reduced it to 20 components and then employed a 3-layer neural network with ReLU activation functions and a mid-layer size of 32. Training data has $4500$ pairwise preferences, and testing data has $500$ pairwise preferences. For gradient-based black-box attacks, we first do PCA to reduce its dimension to size $10$; then we use a 3-layer neural network with mid-layer size $16$.
@@ Appendix B Additional Results
In this section, we present additional results omitted from the main body.
@@
Figure 20 presents the results of trained reward model accuracy in the case of demotion attacks. The results are similar to those for promotion attacks: accuracy degradation is relatively small, demonstrating that the attack is sufficiently stealthy.
Figure 21 presents the ablation results for gradient and RBD methods in the case of demotion attacks.
Finally, Figure 22 shows that the results for multiple target candidates (5, in this case) are comparable to a single target candidate.
@@
Figure 23 presents the results of trained reward model accuracy in the case of demotion attacks in Atari. We can again observe that there is little degradation in training accuracy after the attack.
Figure 24 presents the ablation results for gradient and RBD methods in the case of demotion attacks in the Atari domain.
Figure 25 shows that the results for multiple target candidates are comparable to a single target candidate in the Atari domain.

### v2

@@ 6 Effectiveness of Defense
We study the efficacy of several state-of-the-art defense approaches in the context of poisoning attacks on reward model learning. Specifically, we consider representative defenses from three general classes of approaches (that can be applied in our setting as well) to defend against poisoning attacks: 1) identification and removal of anomalous (poisoned) data, 2) iterative training and removal of high-loss data, and 3) data randomization. As representatives of the first class, we evaluate spectral anomaly detection defense, which arises from the theoretically grounded robust learning literature [84], and the recently proposed Meta-Sift approach [105]. In both spectral outlier and Meta-Sift defense, we first apply the respective approach to identify and remove outliers, and then train the reward model on the rest of the training data. Our representative of the second class is the following three-stage algorithm [21, 52, 87]: 1) train an initial model $R_{\theta_{1}}$, 2) remove $\alpha\cdot B$ datapoints with the largest loss $\mathcal{L}(\tilde{D},\theta_{1})$, where $\alpha\geq 1$ is a hyperparameter (in our experiments, $\alpha=1.5$), and 3) train an updated model on the remaining data. We refer to this as the loss outlier defense. Finally, our representative of the third class of approaches is the ALIBI algorithm that leverages differential privacy and Bayesian post-processing to increase robustness to poisoning attacks [57].
@@
Safety Alignment. First, we evaluate the efficacy of defense approaches in the case of safety alignment for LLM. We exclude Meta-Sift in this setting, which requires iterative training and evaluation of the model, as it is impractical at this scale.
Figure 16: Efficacy of defense against promotion (left) and demotion (right) attacks in the safety alignment (LLM) setting. $B=$ 0.3% of data.
MuJoCo Control. Next, we consider the MuJoCo setting.
Figure 17: Efficacy of defense against promotion attacks (top) and demotion attacks (bottom) in the MuJoCo setting.
The results are shown in Figure 17 for promotion attacks (top) and demotion attacks (bottom). In most cases, we observe that neither the spectral outlier defense nor the loss outlier defense have a significant impact on the attack efficacy. In contrast, ALIBI is considerably more efficacious in both Hopper and Walker environments, although not in the Reacher environment. Nevertheless, with 10% of the data poisoned, attacks remain relatively successful against all of the defenses.
Atari Vision-Based Control. Next, we consider the efficacy of defensive methods against the proposed attacks in the Atari vision-based control domain.
Figure 18: Efficacy of defense against promotion attack (top row) and demotion attacks (bottom row) in the Atari setting.
@@
Recommendation System. In the context of the Amazon recommendation dataset, our results of the relative defense efficacy are shown in Figure 19.
Figure 19: Efficacy of defense against promotion (left) and demotion (right) attacks in the Amazon recommendation setting.
@@
7 Attack Feasibility and Possible Mitigations
Practical exploitation of poisoning attacks remains an open research question. Carlini et al. [9] recently shed light on this by exploiting domain-specific vulnerabilities. In our case of preference poisoning, execution of the attack similarly has to consider how data is collected. To illustrate, the open safety alignment dataset that we use [40], includes $\sim$330K question-answer pairs collected from only 70 annotators. In such cases, even only 4 malicious annotators can poison nearly 6% of the entire dataset.
An important way to mitigate the impact of poisoning attacks explored here is to ensure that datasets are collected from a large number of annotators, with each annotating a relatively small fraction of the data. This would reduce the chances of a small number of malicious labelers significantly subverting reward model training. Similarly, datasets commonly have very few annotators label the same query; we can obtain far greater robustness to poisoning attacks by ensuring that each query has many responses and taking the majority vote.
8 Conclusion
The problem of poisoning attacks generally, and label-flipping in particular, has received much attention in the traditional supervised learning settings, such as classification and regression. Despite some superficial similarity, however, the problem of preference poisoning is structurally distinct, both in terms of the particular optimization problem being solved in learning the reward model, and the particulars of the natural threat model involving promotion and demotion of a target set of outcomes (candidates). Our analysis yields two high-level takeaways with broad significance to security issues surrounding preference elicitation and value alignment. The first is the importance of comprehensive vulnerability analysis that involves a diverse set of attack techniques; in particular, we show that which attack among a collection is best can vary greatly by domain and environment—that is, by the dataset involved. The second is the importance of the particular dataset and data distribution. Indeed, in insightful recent work, [81] observed that the nature of the data distributions (and, thus, datasets) can be essential to robustness of models to data poisoning. This issue, which we highlight empirically, clearly deserves considerably more attention. Moreover, our observation that several state-of-the-art defense techniques fail to provide consistently strong defense in our problem setting suggests that an ability to exert control on the nature of the data we collect can have substantial consequences for the vulnerabilities we have studied here.
Acknowledgments
This work was partially supported by the NSF (IIS-2214141, IIS-1905558, IIS-1903207, CNS-2238635), ARO (W911NF-24-1-0155), DHS (17STQAC00001-06-00), Openphilanthropy, Center for AI Safety for hardware support, and Amazon.
@@ Appendix A Additional Experiment Details
We explored both a simple linear model and a 3-layer neural network with ReLU activation functions and a mid-layer size of 1024. In the PCA scenario, we reduced it to 20 components and then employed a 3-layer neural network with ReLU activation functions and a mid-layer size of 32. Training data has $4500$ pairwise preferences, and testing data has $500$ pairwise preferences. For gradient-based black-box attacks, we first do PCA to reduce its dimension to size $10$; then we use a 3-layer neural network with mid-layer size $16$.
Figure 20: Demotion attack stealth in MuJoCo: test set accuracy.
Figure 21: Relative efficacy of gradient-based demotion attacks (top row) and RBD demotion attacks (bottom row) in MuJoCo.
Figure 22: Attack efficacy in MuJoCo with 5 target candidates.
Figure 23: Demotion attack stealth in Atari: test set accuracy.
Figure 24: Relative efficacy of gradient-based demotion attacks (top row) and RBD demotion attacks (bottom row) in Atari.
Figure 25: Attack efficacy in Atari with 5 target candidates.
Figure 26: Attack efficacy with 5 target candidates on the Amazon recommendation ratings dataset. Neural Network Model (top row) and Linear Model (bottom row).
@@ Appendix B Additional Results
Figure 20 presents the results of trained reward model accuracy for demotion attacks. The results are similar to promotion attacks: accuracy degradation is small, demonstrating that the attack is sufficiently stealthy. Figure 21 presents the ablation results for gradient and RBD methods in the case of demotion attacks. Finally, Figure 22 shows that the results for multiple target candidates (5, in this case) are comparable to a single target candidate.
@@
Figure 23 presents the results of trained reward model accuracy for demotion attacks in Atari. We can again observe that there is little degradation in training accuracy after the attack. Figure 24 presents the ablation results for gradient and RBD methods in the case of demotion attacks in the Atari domain. Figure 25 shows that the results for multiple target candidates are comparable to a single target candidate in the Atari domain.